OpenStudy (anonymous):

-9+|1/4a|=-3 what's the solution of this equation. -3

OpenStudy (anonymous):

choices?

OpenStudy (anonymous):

There is no choices, you're finding the absolute value.

OpenStudy (anonymous):

I am sorry I don't know, do you have any guess at what it could be? If so tell me.

OpenStudy (anonymous):

I worked it down to 9*1/4a=-6*9 But now I'm lost.. Unless the equation is supposed to be 9*1/4a=-6

OpenStudy (anonymous):

You have to keep anything that is in the absolute value bars, in them until they are alone on one side of the = sign

OpenStudy (anonymous):

So it should be something like this -9+|1/4a|=-3 -9 +9 |1/4a|=-3 +9 |1/4a|=6 1/4a*4/1 = 6 * 4/1 a = 24

OpenStudy (anonymous):

so now I would just substitute to make sure I am right -9 +|1/4a|=-3 -9 + |1/4(24)|=-3 -9 + |6| = -3 -9 + 6 = -3 So the answer is a = 24